OpenStudy (anonymous):

solve by graphing 20x+40y=240.00 10x+80y=300.00

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

20x+40y=240.00 10x+80y=300.00 can be simplified as x+2y=1200 x+8y=3000

OpenStudy (anonymous):

okay thank you

OpenStudy (mathmath333):

https://www.desmos.com/calculator/mxhp6yc7ga

OpenStudy (anonymous):

thats what i did but i made 3000 into a decimal and i didn't want that i thought i had to simplify the whole thing

OpenStudy (mathmath333):

notice the intersecting point in the graph that will be the answer

OpenStudy (mathmath333):

error it will be this x+2y=12 x+8y=30

OpenStudy (anonymous):

how did you get 30 what did you divide it by

OpenStudy (mathmath333):

https://www.desmos.com/calculator/fgrbiidhjx

OpenStudy (mathmath333):

2nd equation is this right ? 10x+80y=300

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

divide it by 10

OpenStudy (anonymous):

ooo i thought i had to divide everything by 80y so i can get it by itself

OpenStudy (mathmath333):

in question it is given that u need to solve it by graphing

OpenStudy (mathmath333):

so u have to take 2 points of each equation and graph it on paper of pc whatever ur master told u

OpenStudy (anonymous):

okay thank you